Night Plan UT 30-Dec-2020



Split Night: 1st half LBTI, 2nd half D/E (this plan)

Night Center 00:22

Handover 00:16

Morning Astronomical Twilight: 05:53

NO chance to open Due to Snow and Ice

Nightlog in google sheets here

Night Plan

  • DONE Deploy TCS update after handovoer
  • SOUL AO Checkout following problems of 20201220 UT- CRITICAL - 2hours
    • OSCO at handover - 30min - NCPA on / NCPA off
    • AO Scipts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/AO_Testing/
    • Bin3 Test 0.5-0.75 hr - HIGH priority
      • closed loop data 6 positons of the instrument rotator
      • test the camera lens limits with new M3 position
      • Can use either IDL presets, or AO stars from OT in above directory
    • Bin 1-4 tests on-sky - check functionality
      • use scripts generated from OT in above directory
    • Other Issues:
      • IT 8282 - Bin? pupil files issue
      • IT 8281 - AGW1 loses star at DETXY 10" offset
      • IT 8284 - GOPT issues (pending daytime work too)
      • IT 8271 - preset gains for ESM needed - work on creating the necessary values
    • Observe ESM fields using full AO but with N375 camera - 1 hour per field
      • M67
      • HCG 31
    • Additional?
  • ESM checkout - (M3 and AO checkout must be completed first) - if not started or completed roll over to another night
    • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
    • on-sky checks Bin 2 400Hz (FLAO config) vs Bin 2 100Hz (SOUL config) - 1-2 hrs
    • Collect ESM data - <1" seeing - stable - 1hr minmum
      • Palomar 2 - if clear take photometric standards - probing sensitivity vs FLAO - sets at 3am
  • MODS Checkout - Medium Priority
  • LBC Technical Work - Medium Priority - unlikely to be useful given the HUGE delta-temp for mirrors and ambient
    - O.Kuhn running Technical Work
    • focus offset checks - stable seeing (1" or better on DIMM) - 1.5hr
      • scripts in lbto@robs:~/Service_Observing_D_E_Time_UT_Dec28_Jan02/LBC_R_focus
      • check LBCR R focus, mainly.
        • Focus/collimate, then run superfoc at V+r, followed by a superfoc at [Us,B,g]+R, followed again by superfoc at V+r. Repeat 3x.
    • Gain/Linearity checks - 1hr (photometric conditions, but poor seeing OK)
      • scripts in lbto@robs:~/Service_Observing_D_E_Time_UT_Dec28_Jan02/LBC_Linearity
      • observe a star field with successively longer exposure times, so that a set of stars will have counts that range from 1000 ADU to saturation.
      • several star fields will be chosen, with published photometry (SDSS or other) so we can not only examine the linearity but check the sensitivity.
    • TMS: Run TMS in active mode while collecting a series of pupil images (dfoc = -0.8, Z4/Z11 as sent by fpia) - 2hr any conditions

  • Poor Weather Backups
    • ESM data collection: this is different than ESM tests above - this is sytematic data colleciton - above is checking on paramters for optimal use
      • Scripts: /home/lbto/Service_Observing/D_E_Time_UT_Dec28_Jan_02/ESM_Testing
      • M67 (3 filters and bright and faint AO Ref star)
      • NGC 3921 all filters
    • ICRF3 Spectroscopy (Rothberg)
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/ICRF3_Rothberg_2020B
      • ICR5 93.24-13.51 Highest Priority - cycles through G210 K,H,J, and z (fraternal or mono)
        • sets at 3am
        • Target can sustain 2mags of extinction or more - uses 0.25" slit - J=12, H=9.5, K=7.2
      • Observe telluric first
      • Double Check high Dec objects to make sure telluric/science are close spatially
    • If clear but poor seeing, please observe standard stars, cycling through z,J,H,K,Ks
      • Scripts: /home/lbto/Service_Observing/Nov_Dec_2020_E_D_Nights/Poor_Weather/LUCI_Phot_ZP
      • Select stars to use based on visibility - OT file is in directory, export the scripts as needed
      • Please observe the same star at least two times at different airmasses
-- BarryRothberg - 29 Dec 2020
Topic revision: r4 - 30 Dec 2020, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback